Using one engine of a ferry boat, it takes 6 h longer to cross a channel than it does using a second engine alone. with both engines operating, the ferryboat can make the crossing in 4 h. how long would it take each engine, working alone, to power the ferry boat across the channel?
Answer by TimothyLamb(4379)

x = time for engine A alone
y = time for engine B alone
---
y = x + 6
---
r = w/t
rates are additive:
1/x + 1/y = 1/4
---
1/x + 1/y = 1/4
1/x + 1/(x + 6) = 1/4
1(x + 6)/x(x + 6) + 1x/x(x + 6) = 1/4
1(x + 6) + 1x = (1/4)x(x + 6)
x + 6 + x = (1/4)xx + (6/4)x
(1/4)xx + (6/4)x - 2x - 6 = 0
(1/4)xx + (6/4)x - (8/4)x - 6 = 0
(1/4)xx - (2/4)x - 6 = 0
4*( (1/4)xx - (2/4)x - 6 = 0 )
xx - 2x - 24 = 0
---
the above quadratic equation is in standard form, with a=1, b=-2 and c=-24
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -2 -24
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 6
x = -4
---
the negative root doesn't fit the problem statement, so use the positive root:
---
x = 6
y = x + 6
y = 6 + 6
y = 12
---
answer:
x = time for engine A alone = 6 hours
y = time for engine B alone = 12 hours
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---